OpenStudy (anonymous):

solve for x log x+log(x-1)=log(4x)

OpenStudy (zarkon):

\[\log(x)+\log(x-1)=\log[x(x-1)]\]

OpenStudy (anonymous):

what about the log(4x) ?

OpenStudy (zarkon):

\[\log[x(x-1)]=\log(4x)\Rightarrow x(x-1)=4x\]